OpenStudy (anonymous):

Compare and contrast the two quadratic equations below. In order to receive full credit, use complete sentences to describe the following: The direction each parabola opens The vertex of each parabola y = x2 − 4x y = −2x2 + 8x − 12 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

what did you get

OpenStudy (anonymous):

#1 is up and #2is down, i just dont know the vertex

jimthompson5910 (jim_thompson5910):

both are correct

jimthompson5910 (jim_thompson5910):

to find the vertex, you need to find the x coordinate of the vertex first

jimthompson5910 (jim_thompson5910):

so you use this formula x = -b/(2a)

OpenStudy (anonymous):

Can you help me by setting them up.

jimthompson5910 (jim_thompson5910):

the first equation, what are the values of a, b, c?

OpenStudy (anonymous):

a is x and b is 4, but there isn't a c.

jimthompson5910 (jim_thompson5910):

not really

jimthompson5910 (jim_thompson5910):

y = x^2 − 4x is the same as y = 1x^2 − 4x

jimthompson5910 (jim_thompson5910):

so a = 1 and b = -4, c = 0 x = -b/(2a) x = -(-4)/(2*1) x = 4/2 x = 2

OpenStudy (anonymous):

So, number one is 2?

jimthompson5910 (jim_thompson5910):

so the x coordinate of the vertex is 2 if x = 2, then y = ???

OpenStudy (anonymous):

2?

jimthompson5910 (jim_thompson5910):

y = x^2 − 4x y = (2)^2 − 4(2) y = 4 - 8 y = -4 See how I'm getting this?

OpenStudy (anonymous):

Ohh, yeah. >.< I messed up. Thanks for showing me :)

jimthompson5910 (jim_thompson5910):

sure thing so what this all means is that the vertex of y = x^2 − 4x is (2, -4)